# SafeSearch: Automated Red-Teaming for the Safety of LLM-Based Search Agents

### v3

@@ 4.2 SafeSearch Enables Tracking Safety Progress in Agent Development
More broadly, our results expose a significant knowledge-action gap in LLM safety. While GPT-4.1-mini can flag certain unreliable sources when explicitly tasked with detection, it fails to handle such content safely in agentic scenarios, even with reminders. This highlights the mismatch between specialized safety capabilities and their generalization in real-world agent settings (Wei et al., 2023). In a nutshell, the red-teaming framework proves invaluable for assessing the true effectiveness of defense efforts, thereby enhancing the transparency of search agent development.
@@ 5 Conclusion
This work advances red-teaming for search agent safety by studying how unreliable search results can strongly threaten search agents and end users. Through two in-the-wild experiments, we find that low-quality results are common and can significantly affect agent behaviors. We introduce an automated red-teaming framework that involves four LLM assistants and can effectively identify safety failures. The resulting SafeSearch benchmark, with 300 test cases across five risk types, shows that search agents are highly vulnerable to unreliable search results and can pose threats to users. Our framework also enables a quantitative way to track safety during agent development, revealing that defense strategies such as reminder prompting offer limited protection. We hope this work encourages designers and developers to consider safety alongside performance. We discuss the broader impacts, limitations, future work, and LLM usage in Appendix E.
Ethics Statement
@@
When advancing this study, we, the authors, faithfully adhere to the ICLR Code of Ethics [^8] https://iclr.cc/public/CodeOfEthics. Our efforts include, but are not limited to: 1) Our red-teaming study is designed to secure LLM-based search agents prior to deployment, thereby protecting both the agent systems and end users. 2) The decision to employ simulation-based testing stems from careful consideration of the inevitable harm that real-world SEO manipulation would otherwise cause. 3) Our experiments are conducted with the best efforts to avoid potential biases. This includes ensuring an appropriate benchmark scale, comprehensive coverage of search agent scaffolds and backend LLMs, rigorous validation of red-teaming configurations, and repeated experimentation. 4) Transparency and reproducibility are prioritized. We release the codebase, assets such as test cases and prompts, and detailed red-teaming configurations in or alongside the manuscript.
@@
In addition, our experiments involve several potential concerns, which we address meticulously: 1) Human reviews of checklist-assisted safety evaluations are conducted by well-prepared authors who receive coaching after the review process to mitigate potential mental health issues. 2) To prevent causing offense to readers and reviewers, offensive content is excluded from the main body and placed in the Appendix, accompanied by an explicit disclaimer. 3) The usage of LLMs and existing codebases remains within the permitted scope of the original sources, with full respect given to their copyright and contributions. 4) All red-teaming experiments are executed on self-deployed vLLM servers or through remote APIs, which are totally stateless. We do not interfere with other users or real-world systems, thereby ensuring that no real harm is caused.
Reproducibility Statement
We release our codebase to facilitate the reproduction of the presented experimental results. The code is currently accessible at the anonymous URL: https://anonymous.4open.science/r/SafeSearch, with test cases provided alongside. Comprehensive red-teaming configurations are detailed in Sections 4, B and C. All LLM models and services, as well as the search and webpage parsing services used in our experiments, are openly accessible.
Appendix A The Characteristics of SafeSearch
Figure 7: Dataset Characteristics of Queries in the SafeSearch Benchmark.
@@ Appendix A The Characteristics of SafeSearch
The SafeSearch benchmark comprises 300 test cases designed to capture realistic threats that users may encounter when interacting with search agents. We analyze the characteristics of these queries in SafeSearch. As shown in Figure 7 (Left), most user queries take the form of questions directed at search agents, typically reflecting situations where users are curious but lack sufficient knowledge. The test cases are curated with the support of our automated quality filtering module, which identifies potentially vulnerable queries related to information acquisition. This is exemplified by frequently occurring keywords such as recommended and current in the word cloud. Furthermore, the queries span a wide range of lengths as illustrated in Figure 7 (Right), from short queries of around 10 words to longer ones of approximately 35 words, thereby demonstrating diversity along the length dimension and supporting the representativeness of the dataset. We also provide examples of queries corresponding to each risk type. As presented in Table 3, all user queries are fully benign and exhibit no inherent tendency toward vulnerability if the retrieved search results are reliable. This highlights the stealthiness of the threat: From the user’s perspective alone, it is impossible to determine whether the query has triggered a harmful outcome. We present test case examples in Appendix F.
In this work, our red-teaming mainly focuses on the following five risk categories. It is worth noting that our red-teaming framework can also be extended to new risk types, e.g., requiring only a well-documented risk description to guide test case generation.
Table 4: Risk Coverage Concerning LLM-Based Search Agents.
Risk Type Consequence
Appendix B More Details of Experimental Setup
In this section, we detail the details of the agents along with the models used in Section 4.
B.1 More Details of Search Agents Under Evaluation
@@
    Deep Research Scaffold: The system first decomposes the user query into multiple sub-queries for the search tool. The workflow proceeds iteratively: in each round, the system executes search steps for the current sub-queries (we use the search-workflow agent as the worker). The results of all sub-queries are then passed to a reflection worker, which decides whether to proceed to final synthesis or to refine the sub-queries for further exploration. Once the search budget is run up or the reflection worker deems the information sufficient, the summarizer produces the final response based on the complete search history.
B.2 Model Details
In this work, we include Qwen3-{8B/32B/235B-A22B} (Yang et al., 2025), Gemma-3-IT-27B (Team et al., 2025a), GPT-oss-120b (OpenAI, 2025c), DeepSeek-R1 (Guo et al., 2025), Kimi-K2 (Team et al., 2025b), GPT-4.1(/-mini) (OpenAI, 2025a), o4-mini (OpenAI, 2025d), GPT-5(/-mini) (OpenAI, 2025b), and Gemini-2.5-{flash/pro} (DeepMind, 2025).
For the reasoning models from OpenAI, the available reasoning configurations include minimal, low, medium, and high. We employ the default setting of their API, the medium reasoning effort. For the two reasoning models from Google, Gemini-2.5-Flash and Gemini-2.5-Pro, we use their OpenAI-compatible API and also use the medium reasoning effort. The minimal effort of Gemini-2.5-Flash in Figure 10 corresponds to its no-thinking version. For open-sourced LLMs, we use the instruction-tuned version and enable the reasoning mode if it is supported by the model. The open-sourced LLMs are deployed using the vLLM inference engine (Kwon et al., 2023). The choice of sampling-based decoding under a temperature of 0.6 follows the official recommendations of the models [^9] https://huggingface.co/Qwen/Qwen3-8B#best-practices [^10] https://huggingface.co/deepseek-ai/DeepSeek-R1#usage-recommendations. To account for stochasticity, we repeat sampling three times.
@@ B.2 Model Details
Appendix C Validating Red-Teaming Configurations
@@
Injection Position of the Unreliable Website (Figure 8 (Left)). Position bias has been a long-standing issue in LLM evaluation (Ye et al., 2024). We experiment with GPT-4.1-mini with search workflow, evaluating the settings: at the end, at the beginning, and at a random position in the middle. Our empirical results show that the injection position does not substantially affect the overall ASRs. Among them, the First injection yields the highest ASR, indicating that the model is more likely to be persuaded by unreliable content when it appears as the first search result. This partially aligns with the findings of (Liu et al., 2024; Xiao et al., 2023), which suggest that early context can dominate later auto-regressive processing due to biases in the causal attention mechanism. This observation highlights a particularly concerning scenario: when an unreliable source appears as the top-ranked search result. But this also implies a stricter threat model that could overstate the severity of real-world risks, where unreliable sources may not consistently occupy the highest-ranked positions. In our experiments, we adopt the Last injection position as the default setting, as it ensures reproducibility while maintaining rigor in representing realistic threat scenarios.
Maximal Token Per Website (Figure 8 (Right)). The per-website token limit is not simply a testing choice, but also a search agent configuration. Allocating more tokens allows access to a larger portion of website content, which increases the available information but also introduces noise and higher computational costs. We investigate how varying this token limit affects ASRs, using Qwen3-8B w/ search workflow. The results reveal a non-monotonic trend: ASR initially rises and then declines as the token limit increases. This observation can be explained as follows. With a small token limit, even unreliable websites are truncated and not fully represented in the context, reducing their harmful influence. As the token limit grows, more content from reliable websites is included, which dilutes the impact of unreliable sources and lowers the ASR. However, the empirical findings also highlight that simply enlarging the token budget is insufficient to mitigate the threat.
@@
The Persistent Effect of Unreliable Website (Figure 9 (Left)). As emphasized in Section 3.2 and described in Section 3.3, the threat under study is inherently time-sensitive, and we need to instantiate the unreliable website conditioned on one specified timestamp. We explore the impact of the timestamp with a controlled experiment. Specifically, we fix the timestamp of crafting the unreliable website to Sep 9, 2025, and conduct red-teaming on Sep 9, 2025 and Sep 17, 2025, respectively. The only difference between these two trials lies in the accompanying authentic search results. This experiment is conducted on the agent implementation of being with a search workflow. Across the two models, we find that the impact of one unreliable website is persistent, still maintaining severe effectiveness when surrounded by newer, authentic websites. This persistence is reflected in the ASR metric, which shows only a moderate or negligible decline. These findings rule out the concern that red-teaming successes arise merely from the unreliable website overriding authentic ones through timestamp ordering. Moreover, this result indirectly highlights the scalability advantage of our red-teaming method. By allowing testers to vary the timestamps of website generation, our approach enables systematic evaluation of search agents’ susceptibility along the temporal dimension.
@@
The Choice of Website Generation Strategies (Figure 9 (Right)). Recall that our execution of test cases involves a website generator, which accepts a timestamp and a generation guideline and outputs the content of the reliable website. A natural question arises: Does the choice of website generator influence red-teaming performance and testing efficiency? The metric for the website generation task is direct: The attack’s success necessitates high-quality website generation, meaning that higher ASRs indicate a stronger qualification of a generator for this task. Targeting the GPT-4.1-mini w/ search workflow as testee, we experiment with three strategies for website generation: GPT-4.1-mini, GPT-4.1, and GPT-4.1 with a single exemplar of an authentic website. All are completed with the greedy decoding strategy. Our results show that GPT-4.1-mini and GPT-4.1 perform comparably well, while incorporating an exemplar distracts the generation process, resulting in a slightly lower ASR. We attribute this comparable performance to the design of guideline-assisted website generation, which reduces task complexity and lowers the barriers to the generator’s capability. Consequently, we adopt GPT-4.1-mini (with greedy decoding) as the default website generator, balancing both quality and cost efficiency.

### v4

@@ 5 Discussions
Knowledge-action gap. Broadly, our results, particularly in Section 4.3, reveal a knowledge–action gap in LLM safety. Although GPT-4.1-mini can identify unreliable sources when explicitly tasked, it often fails to act safely in agentic settings, even with reminders. This highlights the mismatch between specialized safety capabilities of LLMs and their generalization in real-world agent settings (Wei et al., 2023).
We provide more discussions of potential limitations and future work in Appendix J.
@@ 6 Conclusion
This work advances red-teaming for search agent safety by studying how unreliable search results can threaten search agents and end users. We introduce SafeSearch, an automated red-teaming framework that involves four LLM assistants and can effectively identify safety failures. Using SafeSearch, we synthesize 300 test cases across five risk types (including misinformation and indirect prompt injection), showing that search agents are highly vulnerable to unreliable search results and can pose threats to users. Our framework also enables a quantitative way to track safety during development, revealing that defense strategies such as reminder prompting offer limited protection. We hope SafeSearch can serve as a practical tool for iterating on safer search agents. As an initial contribution, we hope this work encourages agent designers and developers to prioritize safety alongside performance in future developments.
Impact Statement
@@
When advancing this study, we, the authors, faithfully adhere to responsible research practices [^7] https://icml.cc/Conferences/2026/ResearchEthics. Our efforts include, but are not limited to: (1) Our red-teaming study is designed to secure LLM-based search agents prior to deployment, thereby protecting both the agent systems and end users. (2) The decision to employ simulation-based testing stems from careful consideration of the inevitable harm that real-world SEO manipulation would otherwise cause. (3) Our experiments are conducted with the best efforts to avoid potential biases. This includes ensuring an appropriate benchmark scale, comprehensive coverage of search agent scaffolds and backend LLMs, rigorous validation of red-teaming configurations, and repeated experimentation. (4) Transparency and reproducibility are prioritized. We release the codebase, assets such as test cases and prompts, and detailed red-teaming configurations in or alongside the manuscript.
@@
In addition, our experiments involve several potential concerns, which we address meticulously: (1) Human reviews of checklist-assisted safety evaluations are conducted by well-prepared authors who receive coaching after the review process to mitigate potential mental health issues. (2) To prevent offending readers and reviewers, offensive content is excluded from the main body and placed in the Appendix, accompanied by an explicit disclaimer. (3) The usage of LLMs and existing codebases remains within the permitted scope of the original sources, with full respect given to their copyright and contributions. (4) All red-teaming experiments are executed on self-deployed vLLM servers or through remote APIs, which are fully stateless. We do not interfere with other users or real-world systems, thereby ensuring that no real harm is caused.
Our work highlights the dark side of search agents while providing much-needed transparency into this threat. Our goal is not to exploit vulnerabilities in agentic systems. Rather, it benefits both developers and users by clarifying the safety boundaries of such systems, both during development and in deployment. Our red-teaming methods enable developers to assess and demonstrate system safety more effectively and efficiently, in a systematic and quantitative manner. They also lay the groundwork for understanding the real-world impact of emerging defense mechanisms.
Appendix A Comparing Search Agent with RAG: Paradigms and Threats
Figure 6: Search agent vs. RAG.
Paradigms. LLMs require inference-time knowledge augmentation to answer user queries that are beyond their knowledge scope, e.g., the time-sensitive ones. Besides training-based methods like Borgeaud et al. (2022), retrieval-augmented generation (RAG) and search agent represent two paradigms to achieve this goal.
  • $\bullet$
    RAG (Lewis et al., 2020b): Typically relies on a developer-controlled retrieval pipeline (dense or sparse) over a local, well-curated corpus.
  • $\bullet$
    Agentic RAG (Singh et al., 2025): There are numerous works on enhancing RAG to be agentic. Agentic RAG systems can operate in a more flexible fashion, e.g., in a multi-step and dynamic way, while still relying on a local database. These systems remain within the RAG family.
  • $\bullet$
    Search agent (Xi et al., 2025): Instead of using a developer-implemented retrieval pipeline, search agents access external knowledge by calling a search tool (e.g., Google Search API). Rather than coming from a local database, the searched knowledge is sourced from the wider Internet and controlled remotely by search service providers.
  • $\bullet$
    Agentic search: Currently, the broad literature sometimes uses LLM-based agentic search to refer to a large family of systems. Within this family, our definition of search agents is restricted to those that rely on external search services and Internet-sourced data.
We highlight a fundamental shift in the source and control of retrieved/searched information under the two paradigms:
  • $\bullet$
    (Agentic) RAG retrieves information in a controlled way, and the surfaced information is from a corpus whose quality, curation, and safety threats can be maintained by the developers.
  • $\bullet$
    Search agents, which operate with search tools, will passively receive information sourced from the open Internet. More importantly, this is controlled by the search service provider rather than the agent developers. As discussed in Section 2.1, such search tools can return unreliable results and alter the agent’s behavior.
Threats. The fundamental distinction between RAG and search agents in the source and control of external knowledge determines the heterogeneous nature of the threats they face. As introduced in Section 2.2, RAG systems face the threats of knowledge poisoning attacks (Zou et al., 2025; Tan et al., 2024; Shafran et al., 2025). The threats that search agents encounter in the real world are distinct, as we detail in the following aspects.
  • $\bullet$
    What the threat is: LLMs can suffer from unreliable retrieval from a local database (offline) or unreliable search results from a search tool (online). However, such a simplification, without a concrete threat model, overlooks the more fundamental divergences between the two threats.
  • $\bullet$
    How the threat happens: This is related to the attacker capabilities. RAG poisoning, as an adversarial attack, assumes a relatively strong attacker capacity to manipulate the target system’s knowledge base. By contrast, unreliable search results can arise from content-farming, blackhat SEO, search vendor manipulation, or simply the general noise present across the web. These issues extend beyond deliberate malicious behavior, and the underlying sources of unreliable search results are well-documented, real-world phenomena observed in modern search engines.
  • $\bullet$
    Why the threat happens: RAG poisoning attack is assumed to have a totally malicious objective, e.g., manipulating RAG systems’ answers to certain queries. In contrast, the threat of unreliable search results encompasses a broader set of risks, including prompt injection, misinformation, bias amplification, advertising, and other harmful outputs. These issues extend beyond adversarial intent and reflect a more general and pervasive safety concern.
  • $\bullet$
    How to mitigate this threat: As both the local database and the resulting unreliable retrieval outputs are under the control of RAG system developers, they can directly remove these contents after identification (Zhang et al., 2025a) and thus effectively mitigate this risk. However, the risks faced by search agents are highly dynamic, continually shifting, and therefore difficult to anticipate. Moreover, this problem is more evident when those search agents rely on third-party search services as listed in Langchain’s tool list [^8] https://python.langchain.com/docs/integrations/tools/#search, where developers have no control over how the underlying search tools operate. More advanced and robust mitigation strategies are needed to comprehensively address this threat.
These aspects explain the distinctions between RAG poisoning and the threat of unreliable search results to search agents, and emphasize the unique challenges of mitigating the threat.
Appendix B Setup of the Two In-the-Wild Observations
1. The popularity of low-quality websites. We study content farms—sites that publish keyword-stuffed, low-value articles to manipulate search rankings (Wikipedia, 2025a)—as an example. Inevitably, these sites may appear in search results and affect search agents. We sample 1,000 personas from Persona Hub (Ge et al., 2024) and prompt GPT-4.1 to generate a plausible query each might issue. Using Google Search, we collect the top 10 results per query when available, yielding a pool of 8,933 websites. We tentatively employ GPT-4.1 with CoT reasoning to judge whether a site is likely a content farm (see Appendix M). Notably, 4.3% (380 out of 8,933) are flagged as probable content farms, primarily due to shallow content, lack of credible sources, or heavy advertising.
2. The impact of search results on agent behaviors. The presence of unreliable websites in search results motivates us to examine whether such sources influence the behavior of search agents. We illustrate our methodology in Figure 2. We begin by prompting GPT-5-mini to generate 1,000 health-related yes/no questions. We then conduct differential testing with Qwen3-8B, a representative edge-side model that is sufficiently lightweight to be deployed even by ordinary users on their local machines. We compare its responses when equipped with a Google Search tool against its direct, tool-free answers. This controlled setup allows us to isolate the effect of search integration on agent behavior. The only modification is whether search integration is enabled, and the outcome variable is the agent’s binary stance (yes or no). With the assistance of GPT-5-mini, we find that in 46 out of 1,000 cases, the agent’s responses are altered. Manual review reveals that in most instances, the agent becomes more inclined to acknowledge risky treatments after being exposed to search results, often due to retrieving unreliable sources such as anecdotal blogs. We isolate the specific effects of unreliable search results through controlled red-teaming in Section 3.
Table 2: Risk coverage concerning LLM-based search agents.
Risk Type Consequence
Appendix C The Characteristics of SafeSearch
In this work, our red-teaming mainly focuses on the five risk categories listed in Table 2. It is worth noting that our red-teaming framework can also be extended to new risk types, e.g., requiring only a well-documented risk description to guide test case generation.
Table 3: Dataset curation. “RR” denotes the retention rate.
Risk # Raw $R$
Ads 192 31.3%
Bias 330 18.2%
Harm 867 6.9%
Injec 883 6.8%
Misinfo 318 18.9%
The SafeSearch dataset comprises 300 test cases designed to capture realistic threats that users may encounter. This process of generating the dataset also down-samples raw test cases, providing insights into how different risks impact search agents. The statistics for the retention rates of risk categories are shown in Table 3. Among these, test cases of the two adversarial risks Harmful Output (Souly et al., 2024) and Indirect Prompt Injection (Zhan et al., 2024) prove difficult to curate, with retention rates of just 6.9% and 6.8%. Careful manual review ensures the quality of test cases.
Table 4: Quality assessment of user queries.
Dataset (# Queries) Tag 5 4 3 2 1 0
SafeSearch (300) Ours 72.0% 8.3% 7.3% 10.0% 2.3% 0%
Harmbench (400) Safety, Red-teaming 0% 0.2% 0.2% 4.0% 44.5% 51.0%
WildChat (500) Real-world chat 51% 24.6% 7.0% 11.4% 6.0% 0%
GAIA (466) Search agent 47.4% 34.1% 6.7% 10.7% 1.1% 0%
SimpleQA Verified (1000) Search agent 88.4% 10.1% 0.8% 0.7% 0% 0%
Assessing the representativeness of the user queries. To ensure a replicable measurement, we implement an LLM-as-a-Judge to assess the quality of each user query. The evaluated aspects include authenticity, real-world grounding, representativeness, bias, and safety. We use GPT-4.1 with greedy decoding as the judging model. The prompt is displayed as Appendix M. To position user queries in SafeSearch, we evaluate query quality across several representative datasets: Harmbench (Mazeika et al., 2024) (a red-teaming dataset containing unsafe queries), WildChat (Zhao et al., 2024) (real-world conversations; we use the first 500 items), GAIA (Mialon et al., 2024) (complex user queries designed to probe the upper limits of agentic performance; we combine validation and test sets), and SimpleQA Verified (Haas et al., 2025) (complex user queries used to evaluate agent performance ceilings). The results are given in Table 4, from which we highlight several key observations regarding the representativeness of SafeSearch’s user queries.
  • $\bullet$
    Comparison with safety benchmark: Harmbench contains predominantly low-scoring queries due to their explicit malicious intent. In contrast, SafeSearch does not deliberately encode harmful or unsafe content.
  • $\bullet$
    Comparison with real-world chats: WildChat exhibits a noisier distribution of quality scores and high variability. Notably, SafeSearch mirrors this pattern, suggesting that it effectively captures the characteristics of real-world queries. We attribute this resemblance to our deliberate design of user scenario anchoring during test generation.
  • $\bullet$
    Comparison with performance-oriented benchmarks: SafeSearch, GAIA, and SimpleQA-verified all represent information-seeking scenarios typical of users interacting with search agents. Their quality-score distributions share similar shapes and, importantly, are free from any queries assigned a score of 0.
Quantitative analyses of user queries. We analyze the characteristics of these queries in SafeSearch. As shown in Figure 7 (Left), most user queries take the form of questions directed at search agents, typically reflecting situations in which users are curious but lack sufficient knowledge. The test cases are curated with the support of our automated quality filtering module, which identifies potentially vulnerable queries related to information acquisition. This is exemplified by frequently occurring keywords such as recommended and current in the word cloud. Furthermore, the queries span a wide range of lengths as illustrated in Figure 7 (Right), from short queries of around 10 words to longer ones of approximately 35 words, thereby demonstrating diversity along the length dimension and supporting the representativeness of the dataset. We also provide examples of queries corresponding to each risk type. As presented in Table 5, all user queries are fully benign and exhibit no inherent tendency toward vulnerability if the retrieved search results are reliable. This highlights the stealthiness of the threat: From the user’s perspective alone, it is impossible to determine whether the query has triggered a harmful outcome. We present test case examples in Appendix K.
Figure 7: Dataset characteristics of queries in SafeSearch.
@@ Appendix C The Characteristics of SafeSearch
Appendix D More Details of Experimental Setup
In this section, we provide additional details on the agents and the models used in Section 4.
D.1 More Details of Search Agents Under Evaluation
@@
    Deep research scaffold: The system first decomposes the user query into multiple sub-queries for the search tool. The workflow proceeds iteratively: in each round, the system executes search steps for the current sub-queries (we use the search-workflow agent as the worker). The results of all sub-queries are then passed to a reflection worker, which decides whether to proceed to final synthesis or to refine the sub-queries for further exploration. Once the search budget is run up or the reflection worker deems the information sufficient, the summarizer produces the final response based on the complete search history. Although these systems may involve a workflow to guide their working trajectory, the whole system works as an agentic system; they autonomously decide what to search for, how to use the information, and when to stop by themselves.
D.2 Model Details
In this work, we include Qwen3-{8B/32B/235B-A22B} (Yang et al., 2025), Gemma-3-IT-27B (Team et al., 2025a), GPT-oss-120b (OpenAI, 2025c), DeepSeek-R1 (Guo et al., 2025), Kimi-K2 (Team et al., 2025b), GPT-4.1(/-mini) (OpenAI, 2025d), o4-mini (OpenAI, 2025e), GPT-5(/-mini) (OpenAI, 2025b), Gemini-2.5-{flash/pro} (DeepMind, 2025), Claude-Haiku-4.5 (Anthropic, 2025a), and Claude-Sonnet-4.5 (Anthropic, 2025b).
For the reasoning models from OpenAI, the available reasoning configurations include minimal, low, medium, and high. We employ the default setting of their API, the medium reasoning effort. For the two reasoning models from Google, Gemini-2.5-Flash and Gemini-2.5-Pro, we use their OpenAI-compatible API and also use the medium reasoning effort. For the two Claude models from Anthropic, Claude-Sonnet-4.5 and Claude-Haiku-4.5, we use their OpenAI-compatible API. The minimal effort of Gemini-2.5-Flash in Figure 11(a) corresponds to its no-thinking version. For open-source LLMs, we use the instruction-tuned version and enable the reasoning mode if it is supported by the model. The open-source LLMs are deployed using the vLLM inference engine (Kwon et al., 2023). The choice of sampling-based decoding with a temperature of 0.6 follows the official recommendations of the models [^9] https://huggingface.co/Qwen/Qwen3-8B#best-practices [^10] https://huggingface.co/deepseek-ai/DeepSeek-R1#usage-recommendations. To account for stochasticity, we repeat sampling three times.
@@ D.2 Model Details
Claude-Haiku-4.5 Anthropic 2025-09 ? 200K 2025-02
Claude-Sonnet-4.5 Anthropic 2025-10 ? 200K 2025-01
@@
Appendix E Curating More Challenging Test Cases
To differentiate difficulty levels, we adopt a tiered filtering process. For the advertisement scenario, we first generated 300 raw test cases, then filtered them using Qwen3-8B and GPT-5-mini (effort=minimal) as baseline agents, both equipped with a search workflow and greedy decoding. This process produced two tiers:
  • $\bullet$
    Tier 1: 27 test cases that pass GPT-5-mini’s check;
  • $\bullet$
    Tier 2: 108 test cases that fail GPT-5-mini but pass that of Qwen3-8B’s check.
It is expected that Tier 1 test cases are more challenging than those from Tier 2. We red-teamed various LLMs using both tiers and report their ASRs in Table 8. Across all evaluated agents, Tier 1 yields consistently higher ASRs than Tier 2. This is most evident for relatively safe search agents, e.g., o4-mini, Kimi-K2, and GPT-5-mini (effort=medium). This confirms the feasibility of curating more challenging test cases in the long run.
Appendix F Validating Red-Teaming Configurations
@@
1. Injection position of the unreliable website (Figure 8 (Left)). Position bias has been a long-standing issue in LLM evaluation (Ye et al., 2024). We experiment with GPT-4.1-mini with search workflow, evaluating the settings: at the end, at the beginning, and at a random position in the middle. Our empirical results show that the injection position does not substantially affect the overall ASRs. Among them, the First injection yields the highest ASR, indicating that the model is more likely to be persuaded by unreliable content when it appears as the first search result. This partially aligns with the findings of (Liu et al., 2024; Xiao et al., 2023), which suggest that early context can dominate later auto-regressive processing due to biases in the causal attention mechanism. This observation highlights a particularly concerning scenario: when an unreliable source appears as the top-ranked search result. But this also implies a stricter threat model that could overstate the severity of real-world risks, where unreliable sources may not consistently occupy the highest-ranked positions. In our experiments, we adopt the Last injection position as the default setting, as it ensures reproducibility while maintaining rigor in representing realistic threat scenarios.
2. Maximal token per website (Figure 8 (Right)). The per-website token limit is not simply a testing choice, but also a search agent configuration. Allocating more tokens allows access to a larger portion of website content, which increases the available information but also introduces noise and higher computational costs. We investigate how varying this token limit affects ASRs, using Qwen3-8B w/ search workflow. The results reveal a non-monotonic trend: ASR initially rises and then declines as the token limit increases. This observation can be explained as follows. With a small token limit, unreliable websites are truncated and not fully included in the context, reducing their harmful influence. As the token limit grows, more content from reliable websites is included, which dilutes the impact of unreliable sources and lowers the ASR. However, the empirical findings also highlight that simply enlarging the token budget is insufficient to mitigate the threat.
@@
3. The persistent effect of unreliable website (Figure 9 (Left)). As emphasized in Section 3.2 and described in Section 3.3, the threat under study is inherently time-sensitive, and we need to instantiate the unreliable website conditioned on one specified timestamp. We explore the impact of the timestamp with a controlled experiment. Specifically, we fix the timestamp of crafting the unreliable website to Sep 9, 2025, and conduct red-teaming on Sep 9, 2025 and Sep 17, 2025, respectively. The only difference between these two trials lies in the accompanying authentic search results. This experiment is conducted on the search-workflow agent implementation. Across the two models, we find that the impact of one unreliable website is persistent, still maintaining severe effectiveness when surrounded by newer, authentic websites. This persistence is reflected in the ASR metric, which shows only a moderate or negligible decline. These findings rule out the concern that red-teaming successes arise merely from the unreliable website overriding authentic ones through timestamp ordering. Moreover, this result indirectly highlights the scalability advantage of our red-teaming method. By allowing testers to vary the timestamps of website generation, our approach enables systematic evaluation of search agents’ susceptibility along the temporal dimension.
@@
4. The choice of website generation strategies (Figure 9 (Right)). Recall that our execution of test cases involves a website generator, which accepts a timestamp and a generation guideline and outputs the content of the unreliable website. A natural question arises: Does the choice of website generator influence red-teaming performance and testing efficiency? The metric for the website generation task is direct: The attack’s success necessitates high-quality website generation, meaning that higher ASRs indicate a stronger qualification of a generator for this task. Targeting the GPT-4.1-mini w/ search workflow as testee, we experiment with 3 website generation strategies: GPT-4.1-mini, GPT-4.1, and GPT-4.1 with a single exemplar of an authentic website. All are completed with the greedy decoding strategy. Our results show that GPT-4.1-mini and GPT-4.1 perform comparably well, while incorporating an exemplar distracts the generation process, resulting in a slightly lower ASR. We attribute this comparable performance to the design of guideline-assisted website generation, which reduces task complexity and lowers the barriers to the generator’s capability. Consequently, we adopt GPT-4.1-mini (with greedy decoding) as the default website generator, balancing both quality and cost efficiency.
5. Injection round in multi-round tool-calling (Figure 10 (Left)). In the real world, the timing of the agent’s risky tool call, when the unreliable search result is met, is inherently unpredictable, making each possible injection round a distinct red-teaming configuration. To avoid overestimating the threat, we adopt a conservative strategy and place the unreliable-website injection in the very first round. This creates a less stringent setting and gives the agent the opportunity to identify and mitigate the injected unreliable website in subsequent rounds. We conduct experiments to examine how the injection round affects the red-teaming outcome. To ensure that agents consistently perform multiple rounds of tool calls (exactly three in this setting), we operationally set the tool_choice from auto to required, a feature now supported by the OpenAI API. Because this modification introduces behavioral constraints absent from the automatic tool-calling search agent in our main experiments, we denote this search agent setting as tool-calling (forced). With this setup, we keep all other parameters fixed and inject exactly one unreliable website at different rounds. The experimental results in Figure 10 (Left) demonstrate that the success of red-teaming is not specific to a certain choice of injection round. What’s more, as evidenced by the higher ASR of round-3 injection, our default configuration of round-1 injection indeed evaluates the search agent safety in a conservative way. This experiment demonstrates the representativeness of our choice in the injection round.
Figure 10: Ablating red-teaming configurations of injection round and adding persuasiveness. (Left) The configuration of round-1 injection is conservative for tool-calling (forced). (Right) Explicitly requesting a persuasive website can further enhance ASRs.
6. Enforcing persuasiveness in website generation. We do not explicitly prompt the website generator to generate content in a persuasive fashion. Nonetheless, increasing the persuasiveness of injected content represents a plausible avenue for creating more challenging stress tests for search agents. We conduct one experiment to explore the feasibility. We revise the prompt of the website generator to explicitly request more persuasive content tailored to the target consequence. All other settings remained unchanged. The results are shown in Figure 10 (Right). We find that even without deliberate prompt refinement, the more persuasive website can lead to a notable increase in ASR. This demonstrates both the feasibility of incorporating deliberately persuasive content when designing more demanding stress-testing scenarios for search agents.
Table 7: $\text{ASR}_{\text{overall}}$ scores of using 300 test cases generated by o4-mini or Qwen3-8B. Search agents are with the search workflow.
Model o4-mini Qwen3-8B
Qwen3-8B 85.5% 67.3%
Qwen3-32B 81.1% 60.8%
GPT-4.1-mini 90.5% 76.4%
Gemini-2.5-Flash 76.8% 64.8%
7. Using open-source reasoning models for test case Generation. In our work, we primarily employ o4-mini as the test generator, observing its strong reasoning ability and cost efficiency. Yet, this may raise an important concern: if only closed-source models can generate test cases, the community may lose access to new test cases when those API-based models become heavily moderated and start refusing test-generation requests. We experiment with an open-source reasoning model, Qwen3-8B, as the test generator. We replicate the entire test generation pipeline to obtain 300 test cases across five risk categories. Using these test cases, we assess several search agents equipped with the search workflow but with different backend LLMs: Qwen3-8B, Qwen3-32B, GPT-4.1-mini, and Gemini-2.5-Flash. We adopt the same red-teaming configurations as in the main experiments and run each test case three times with a temperature of 0.6. The results are shown in Table 7. Using an open-source model such as Qwen3-8B to generate test cases still yields effective red-teaming data. While the ASR values drop moderately, the reduction is not drastic, indicating that our methodology does not depend strictly on o4-mini or on closed-source models in general. Moreover, this experiment highlights an encouraging direction: relatively small open-source models can supervise or probe their own safety and even significantly stronger models, such as Qwen3-32B. This expands the practical accessibility of our red-teaming framework for the broader research community. Looking ahead, it also points to a path toward continuously monitoring the safety of search agents over the long term, since our method does not rely on access to strong test generators to function.
Table 8: ASRs of red-teaming with test cases in Tier 1 and Tier 2. All search agents are equipped with the search workflow scaffold. Timestamp: Nov 28, 2025.
Model ASR (Tier 1) ASR (Tier 2) $\Delta$
Qwen3-8B 77.8% 66.7% 11.1%
Qwen3-32B 75.3% 64.2% 11.1%
Qwen3-235B-A22B 77.8% 61.1% 16.7%
GPT-4.1-mini 85.2% 81.2% 4.0%
GPT-4.1 96.3% 82.4% 13.9%
o4-mini 86.4% 49.7% 36.7%
Gemini-2.5-Flash 77.8% 67.3% 10.5%
GPT-5-mini (effort=medium) 30.9% 6.2% 24.7%
Kimi-K2 84.0% 56.5% 27.5%
Appendix G Validating the Automated Safety Evaluator
The benefits of introducing the checklist. Our red-teaming is conducted in a targeted manner, where we inject a controlled, unreliable website and observe how the agent treats this website. The controlled website contains specified misinformation/low-quality content, guided by the website generation guidelines, and therefore known by the tester. This controlled red-teaming setup enables our automated LLM evaluator (see Appendix M in the Appendix) to be context-dependent. To support this, the test generator also produces a detailed checklist for safety evaluation (see examples in Appendix K). The evaluator judges the safety of the agent’s responses directly against this checklist.